For a normal eye, the focal length of the lens if f = 1.7 cm. If your friend is nearsighted such that the focal length of her lens is 2.0 cm. What should be the focal length of the corrective lens?
Answer #1

1/f = 1/v + 1/u where f is the focal length of the lens, v is the image distance, and u is the object distance. Because the image is a virtual one, v is given a negative sign

1/f = 1/1.7-1/2

f = 11.33 cm